OpenStudy (anonymous):

An experiment has three outcomes. I, II, III. If ouctcomes I is twice as likely as outcome II, and outcome II is three times as likely as outcome III, what are the probability value of the three outcomes?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (sonyalee77):

P(event I) = 6x P(event II) = 3x P(event III) = x Since these are the only outcomes then P(event I) + P(event II) + P(event III) = 1 10x = 1 so x = 1/10 3x = 3/10 6x = 6/10 = 3/5 so P(event I) = 3/5; P(event II) = 3/10; P(event III) = 1/10

OpenStudy (sonyalee77):

is this good?